As of 2:45pm the CS website appears to be fully restored to service. CIS is still reporting that service outages continue, so we will continue to monitor the situation.
Archives October 2017
Due to campus-wide service outages, the main sections of the CS Dept website, including the home page, have been inaccessible for an extended period of time today, and remain down. We are monitoring CIS' progress in addressing their issues, and we will ensure that the CS website returns to service when they are resolved.
On Wednesday, Oct. 18th, between 7am and 8am, the Networking Group will be performing a VLAN migration of the Self Managed Newtork (10.116.70.0/24) from vlan 4007 to vlan 3962. There will be a brief network connectivity problem during the maintenance windows for user servers, desktops and laptops on the 10.116 ...